15.1 A 39.8 mL sample of a 0.448 M aqueous hypochlorous acid solution is titrated with...

15.1
A 39.8 mL sample of a 0.448 M aqueous hypochlorous acid solution is titrated with a 0.234 M aqueous potassium hydroxide solution. What is the pH after 50.5 mL of base have been added?

15.2

What is the pH at the equivalence point in the titration of a 24.9 mL sample of a 0.313 M aqueous hypochlorous acid solution with a 0.479 M aqueous barium hydroxide solution?

12.1

How many grams of solid ammonium bromide should be added to 1.00 L of a 0.101 M ammonia solution to prepare a buffer with a pH of 9.990 ?

12.2

How many grams of solid sodium cyanide should be added to 1.00 L of a 0.132 M hydrocyanic acid solution to prepare a buffer with a pH of 8.401 ?
Answer #1

is. I ... pla of Hocd = 7:53 initial moles of Holl = 39.8 ML X 0.448m = 39.8x163x *0.448 molly = 17-8304x16%mo? Imitial molesHou and Kou react in til ratio . moles of kod tormed = moles of KOH moc = 11.817x1ó? mol mole of Hod left in solution nuoa =pH = 7.53+0.2934 = 7.8234 15.? At the equivalence point all Hod is converted to · Roch . .. moles of kod = initial moles of Hconci of kold at . molel of kod equivalence point - Total volume = 7.7937X163 mol 33.035 mL. . . koci].0.23592.m pkb of kollNeglect a in denominator because it is small 3:38x167 = x2 => x=023592x3:38 x1o7 . 0.23592 32= 07.974 x16 8 =) *= 17.974 x104plate of NH3 = 4.75 , p. 14-24=14-9.99 Henderson Hamalbalch equation in poM = pkb + log moles of NHA BV moles of NH3 46 4.5 4magn of NH BV should be added = 0.018379 mol x molar mash of NHq BV = 0.018379 mol x 97.94 g/mol @ 1.8 ġ 12.2 pra pka + log m